OpenStudy (baby456):

the regular price for a video game is 164.50 but is on sale for 30% off. What is the amount of the discount? What is the sale price? help( for a medal)!

OpenStudy (anonymous):

discount amount = 164.50 * 0.3 sale price = 164.50 * 0.7